OpenStudy (anonymous):

the bus and train left the same town and headed south at 8 a.m. at noon, the bus was 100 miles behind the train. how far did each one travel if the speed of the train was twice the speed of the bus? the bus and train left the same town and headed south at 8 a.m. at noon, the bus was 100 miles behind the train. how far did each one travel if the speed of the train was twice the speed of the bus? @Mathematics

OpenStudy (anonymous):

r t = d Let r be the speed of the bus. The speed of the train is 2r from the problem statement. Solve the following for r and d: { r* 4 = d, (2*r) 4 = d + 100 } r = 25 mph and d = 100 miles Bus: 25+4 = 100 miles train: 2*25*4 = 200 miles